OpenStudy (anonymous):

What is the equation of a parabola with vertex at (1, −3) and focus at (−2, −3)?

OpenStudy (anonymous):

(x+2)^2 + (y+3)^2 = (x-4)^2

OpenStudy (anonymous):

View Lab The coordinates of the vertex of the parabola 2(y − 1)2 = x + 2 are